It takes Rob 3 hours to do a job. It takes Tim 5 hours to do the same job. If Rob and Tim worked together, how long would it take them? Please advise. Thank you.

Answer by nerdybill(7384)

It takes Rob 3 hours to do a job. It takes Tim 5 hours to do the same job. If Rob and Tim worked together, how long would it take them?
.
Let x = time it would take for them to working together
then
x(1/3 + 1/5) = 1
multiplying both sides by 15:
x(5 + 3) = 15
8x = 15
x = 15/8
x = 1.875 hours
or
x = 1 hour and 53 minutes